OpenStudy (anonymous):

When expanding ln to a fraction, do you only distribute it to the numerator?

OpenStudy (unklerhaukus):

\[a\times\frac{b}{c}=\frac{ab}c\]

OpenStudy (anonymous):

Okay, great, thank you!

OpenStudy (unklerhaukus):

or you "could" go \[a\times\frac{b}{c}=\frac{b}{c/a}\]

OpenStudy (anonymous):

In my problem I have to expand the problem ln(2x^6+3squareroot(x+8)/(x-1)^4

OpenStudy (anonymous):

So just distributing to the numerator, assuming ita ln/1 would prbably be best, I think?

OpenStudy (unklerhaukus):

where is the closing bracket on the natural log?

OpenStudy (anonymous):

It is supposed to be applied to the whole fraction.... sorry if it was confusing!

OpenStudy (unklerhaukus):

\[\ln\left(\frac{2x^6+3\sqrt{(x+8)}}{(x-1)^4}\right)\]?

OpenStudy (anonymous):

Yes!

OpenStudy (unklerhaukus):

well the thing about logs ... \[\log ab=\log a+\log b\] \[\log \frac ab=\log a-\log b\]

OpenStudy (anonymous):

So would I turn the numerator into ln((2x^6)(3sq.rt.x+8))?

OpenStudy (unklerhaukus):

where'd the plus go? \[\log a+b≠\log a+\log b=\log ab\]

OpenStudy (anonymous):

Oh! I missunderstood. So now I have \[\ln \left( 2x ^{6} \right)+ \ln \left( 3\sqrt{x+8} \right)\]

OpenStudy (anonymous):

In the numerator.

OpenStudy (anonymous):

Thanks for all your help!

OpenStudy (unklerhaukus):

no you cant do that the whole numerator will stay in a single log, however you can simplify the log of the denominator